OpenStudy (anonymous):

Solve the following system. -3x + 2y + z = -6 x + 3y + 2z = 5 4x + 4y + 3z = 13. I got z= -5, y=4 and x=3. My teacher told me I did not show my work and she didn't wouldn't take credit for it, even if it was right. Please step out the problems for me and show me what I need to do to get my answer, how to show my work and solve for it. I really need to know this question. SOMEONE please help. I'm desperate.

OpenStudy (texaschic101):

-3x + 2y + z = - 6 ---->(-2) x + 3y + 2z = 5 ---------------- 6x - 4y - 2z = 12 (result of multiplying by -2) x + 3y + 2z = 5 ----------------add 7x - y = 17 what you eliminate in the first 2 equations, you must eliminate in the second 2. x + 3y + 2z = 5 ---->(3) 4x + 4y + 3z = 13 --->(-2) -------------- 3x + 9y + 6z = 15 (result of multiplying by 3) -8x - 8y - 6z = - 26 (result of multiplying by -2) --------------add -5x + y = - 11 now take your results 7x - y = 17 -5x + y = -11 -------------add 2x = 6 x = 6/2 x = 3 now we sub into an equation that does not have a z in it. -5x + y = -11 -5(3) + y = -11 -15 + y = -11 y = -11 + 15 y = 4 now sub back into original equation x + 3y + 2z = 5 3 + 3(4) + 2z = 5 3 + 12 + 2z = 5 2z = 5 - 15 2z = - 10 z = -10/2 z = -5 now we check.. -3x + 2y + z = -6 -3(3) + 2(4) + (-5) = -6 -9 + 8 - 5 = -6 -14 + 8 = -6 -6 = -6 (correct) x = 3, y = 4, and z = -5

OpenStudy (texaschic101):

anytime...its not really that hard, just a little time consuming